OpenStudy (anonymous):

Simplify :) 1 over cube root of 1/x^2 plus cube root of 1/x plus 1

OpenStudy (lgbasallote):

\(\LARGE \frac{1}{\sqrt[3]{\frac{1}{x^2}}} + \sqrt[3]{\frac{1}{x}} + 1\) is the question?

OpenStudy (anonymous):

\[1\over \sqrt[3]{1\over x^2} + \sqrt[3]{1 \over x} + 1\]